Half the difference between two numbers is 2. The sum of the greater number and twice the smaller number is 3. Find the numbers.
Answer by stanbon(75887)

Half the difference between two numbers is 2. The sum of the greater number and twice the smaller number is 3. Find the numbers.
-----
(1/2)(x-y) = 2
x + 2y = 3
-----
Modify:
x - y = 4
x + 2y= 3
-----
subtract and solve for "y":
3y = -1
y = -1/3
----
Solve for "x":
x = y+4
x = 3 2/3
=================
Cheers,
Stan H.
===============